OpenStudy (anonymous):

Which number is equal to 3^3? 7 25 27 66

OpenStudy (anonymous):

3^3 = 3*3*3 so 3*3 = 9 then multiply 9 by 3 and that's your answer

OpenStudy (anonymous):

@IanK16. @Techno openstudy.com/code-of-conduct